# A Survey of Low-bit Large Language Models: Basics, Systems, and Algorithms

### v2

@@ 5.3.2 Evaluation
The evaluation in the benchmarks showcases the most interesting aspects of quantization LLMs, i.e., efficiency and generation quality. We list the detailed tracks in Table 5. For efficiency, the inference efficiency is measured by deployability and throughput, which are the most crucial features in LLMs compression Lin et al. (2024b); Gong et al. (2024). Typically, reducing the storage of the parameters can speed up the inference theoretically, but it depends on the actual system implementation. The benchmark provides us with a fair and convenient probe to distinguish the algorithms and implementations that have practical acceleration and storage saving. The production efficiency is measured by calibration time, which indicates the time and computational resources cost of the PTQ algorithms Gong et al. (2024). Methods that spare lots of resources usually have better generation quality, while those that require less time may have worse generation performance. It is a trade-off in producing quantized LLMs. For generation quality, it has many aspects, such as perplexity, accuracy, logic, completion, trustworthiness and so on Lin et al. (2024b); Gong et al. (2024); Li et al. (2024c); Liu et al. (2024d). Most benchmarks evaluate emergent capability, which is the key feature of LLMs. Specifically, models and algorithms are tested under diverse scenarios, like dialogue, long-context, or multi-task Li et al. (2024c). And some benchmarks are aware of the safety of generative contents, and estimate the trustworthiness and robustness of LLMs Li et al. (2024c); Liu et al. (2024d).
@@ 6 Future Trends and Directions
Quantization Techniques. Despite progress, several challenges remain in quantization techniques. Firstly, one major issue is the unclear origin of outliers in large language models (LLMs), which presents a significant barrier to further reducing quantization bit widths. Research aimed at uncovering the internal mechanisms behind these outliers is crucial and will provide valuable insights for the community, potentially advancing the state of quantization and enabling more efficient models. Secondly, pushing the boundaries of minimal bit representation with acceptable accuracy is highly valuable. Achieving the lowest possible bit width while maintaining performance can fully leverage hardware capabilities and maximize its potential. Thirdly, exploring unified strategies for mixed-bit quantization, including both bit selection and intra-layer/inter-layer bit allocation, is essential for optimizing model performance and efficiency. Current methods primarily emphasize intra-layer mixed precision, often overlooking the potential benefits of inter-layer mixed precision. Last but not least, developing semantic-guided strategies for achieving even lower-bit quantization and compression of key-value (KV) caches will be a major focus. During inference with long context lengths, the primary bottleneck often lies in the substantial memory usage of KV caches. Therefore, identifying effective methods for compressing KV caches is crucial for overcoming this limitation and enhancing model efficiency.

### v3

@@ 5.3.2 Evaluation
The evaluation in the benchmarks showcases the most interesting aspects of quantization LLMs, i.e., efficiency and generation quality. We list the detailed tracks in Table 5. For efficiency, the inference efficiency is measured by deployability and throughput, which are the most crucial features in LLMs compression (lin2024qserve; gong2024llm). Typically, reducing the storage of the parameters can speed up the inference theoretically, but it depends on the actual system implementation. The benchmark provides us with a fair and convenient probe to distinguish the algorithms and implementations that have practical acceleration and storage saving. The production efficiency is measured by calibration time, which indicates the time and computational resources cost of the PTQ algorithms (gong2024llm). Methods that spare lots of resources usually have better generation quality, while those that require less time may have worse generation performance. It is a trade-off in producing quantized LLMs. For generation quality, it has many aspects, such as perplexity, accuracy, logic, completion, trustworthiness and so on (lin2024qserve; gong2024llm; li2024evaluating; liu2024evaluating). Most benchmarks evaluate emergent capability, which is the key feature of LLMs. Specifically, models and algorithms are tested under diverse scenarios, like dialogue, long-context, or multi-task (li2024evaluating). And some benchmarks are aware of the safety of generative contents, and estimate the trustworthiness and robustness of LLMs (li2024evaluating; liu2024evaluating).
Takeaways of subsection 5.3 If your goal is to reproduce a variety of quantization algorithms, LLMC, MI-optimize, and LMQuant are recommended, as they provide a comprehensive suite of quantization methods. If your focus is to deploy across inference frameworks, LLMC stands out as an ideal choice, providing flexible quantization settings and seamless compatibility with multiple backends.
@@ 6 Future Trends and Directions
Quantization Techniques. Despite progress, several challenges remain in quantization techniques. Firstly, one major issue is the unclear low-semantic-valueorigin of outliers in large language models (LLMs), which presents a significant barrier to further reducing quantization bit widths. Research aimed at uncovering the internal mechanisms behind these outliers is crucial and will provide valuable insights for the community, potentially advancing the state of quantization and enabling more efficient models. Secondly, pushing the boundaries of minimal bit representation with acceptable accuracy is highly valuable. Achieving the lowest possible bit width while maintaining performance can fully leverage hardware capabilities and maximize its potential. Thirdly, exploring unified strategies for mixed-bit quantization, including both bit selection and intra-layer/inter-layer bit allocation, is essential for optimizing model performance and efficiency. Current methods primarily emphasize intra-layer mixed precision, often overlooking the potential benefits of inter-layer mixed precision. Last but not least, developing semantic-guided strategies for achieving even lower-bit quantization and compression of key-value (KV) caches will be a major focus. During inference with long context lengths, the primary bottleneck often lies in the substantial memory usage of KV caches. Therefore, identifying effective methods for compressing KV caches is crucial for overcoming this limitation and enhancing model efficiency.